# Explicit Second-Order Min-Max Optimization: Practical Algorithms and Complexity Analysis

### v5

@@ 4.4 Finite-sum min-max optimization
Compared to Lemma 4.9, the computation of sampling probability in Lemma 4.11 requires going through the whole dataset and the cost is $O((m+n)N)$. Nonetheless, the computational savings with smaller sample size dominates such extra cost of computing the sampling probability in a minimization setting (Xu et al., 2016). In particular, the sample size from Lemma 4.11 is smaller as $B_{\textnormal{avg}}\ll B_{\max}$ which occurs if one $B_{i}$ is much larger than the others. In addition, the sample size is proportional to the log of the failure probability in Lemma 4.9 and 4.11, allowing the use of a very small failure probability without increasing the sample size significantly.
@@
To guarantee an overall accumulative success probability of $1-\delta$ across all $T$ iterations, it suffices to set the per-iteration failure probability as $1-\sqrt[T]{1-\delta}$ as we have done in Algorithm 3. Moreover, we have $1-\sqrt[T]{1-\delta}=O(\frac{\delta}{T})=O(\delta\epsilon^{2/3})$. Since this failure probability has only been proven to appear in the logarithmic factor for the sample size in both Lemma 4.9 and 4.11, the extra cost will not be dominating. As such, when Algorithm 3 terminates, all of the Jacobian approximations have satisfied Eq. (4.18). This completes the proof. $\square$

### v6

@@ 4.4 Finite-sum min-max optimization
Compared to Lemma 4.11, the computation of sampling probability in Lemma 4.13 requires going through the whole dataset and the cost is $O((m+n)N)$. Nonetheless, the computational savings with smaller sample size dominates such extra cost of computing the sampling probability in a minimization setting (Xu et al., 2016). In particular, the sample size from Lemma 4.13 is smaller as $B_{\textnormal{avg}}\ll B_{\max}$ which occurs if one $B_{i}$ is much larger than the others. In addition, the sample size is proportional to the log of the failure probability in Lemma 4.11 and 4.13, allowing the use of a very small failure probability without increasing the sample size significantly.
@@
To guarantee an overall accumulative success probability of $1-\delta$ across all $T$ iterations, it suffices to set the per-iteration failure probability as $1-\sqrt[T]{1-\delta}$ as we have done in Algorithm 3. Moreover, we have $1-\sqrt[T]{1-\delta}=O(\frac{\delta}{T})=O(\delta\epsilon^{2/3})$. Since this failure probability has only been proven to appear in the logarithmic factor for the sample size in both Lemma 4.11 and 4.13, the extra cost will not be dominating. As such, when Algorithm 3 terminates, all of the Jacobian approximations have satisfied Eq. (4.18). This completes the proof. $\square$